# 1]organization=Center for Research and Formation in Artificial Intelligence, Universidad de los Andes, Colombia, addressline=Carrera 1 No. 18a-12, city=Bogotá, postcode=111711, country=Colombia

### v1

@@ 5.2 Gene Prediction Benchmark
Figure 6 shows the performance of all methods for every dataset when trained under our two scenarios (raw and SpaCKLE-completed data). It is clear that the prediction performance significantly improves when applying SpaCKLE to every dataset and, in some cases, the best PCC increases to 0.36 points (AHSCC). This result pinpoints the importance of acknowledging missing data for the prediction task and proves the significance of including gene completion in ST pipelines.
@@
Comparing datasets’ difficulty, we find that the most challenging dataset to predict was MMBR (PCC=0.16), while 10XGMBSP emerged as the least difficult (PCC=0.74). When inspecting each dataset’s characteristics, we observe that the organism does not appear to have a significant impact on the difficulty of the task, as the mean prediction PCC achieved for mouse and human datasets is very similar. Furthermore, a larger number of available genes (due to better quality) facilitates prediction, which is evident in a higher average and maximum performance on the datasets with 128 genes compared to those with 32 genes. Finally, results also demonstrate that generalizing in an intra-patient manner typically makes the prediction easier than inter-patient (See Fig.7. a).
@@
We analyze the prediction performance across various tissue types. Figure 7. b presents the performance of each dataset in SpaRED, categorized by tissue type, where the bars indicate the average PCC for each type of tissue. On average, the best prediction performance was observed for skin tissue, while the lowest was for kidney tissue. However, the distribution of tissue types in SpaRED is highly imbalanced, with some tissue types being underrepresented. Notably, both skin and kidney tissues are represented by only a single dataset, making it unreliable to draw definitive conclusions about whether certain tissues are inherently easier to predict than others. The observed differences may be influenced by dataset-specific characteristics rather than general tissue properties.
@@
Hist2ST HGGEP ShuffleNet STNet EGN EGGN HisToGene ViT-B VIT-B+$\Delta$ SEPAL BLEEP
@@
We display the results of evaluating the 8 state-of-the-art models on SpaRED, as well as the baseline experiments on Fig. 8.a sorted by best average performance. The normalized MSE metric indicates how close every model’s results are to the best performance achieved on each dataset. Results show that ViT-B+$\Delta$ attains the best gene expression predictions on average, despite being one of the most straightforward approaches for the prediction task. Moreover, the pie chart showcases that STNet and ViT-B+$\Delta$ emerge most frequently as the best methods. Interestingly, we notice that SEPAL, which is built on top of ViT-B+$\Delta$, falls behind the latter. This contrast reveals that incorporating local vicinity information does not necessarily improve the outputs and that focusing on predicting the $\Delta$ from the mean expression is already a powerful strategy.
Table 3 illustrates the statistical differences in MSE performance across all datasets. A Dunn test with a 5% significance level reveals that most methods do not exhibit statistically significant differences in performance. While previous results indicate that ViT-B+$\Delta$ and STNet most frequently achieve the best results, with ViT-B+$\Delta$ obtaining the highest average performance, the statistical analysis confirms that these improvements are not significant when compared to most other methods. Notably, ViT-B+$\Delta$ shows a statistically significant advantage only over Hist2ST, HGGEP, HisToGene, and BLEEP, while no significant advantage is observed over the other methods. These results indicate that no single state-of-the-art method is definitively superior, highlighting that the existing strategies for improving gene expression prediction remain insufficient. This underscores the need for novel approaches to enhance ST-related tasks.
@@ Acknowledgments
Gabriel Mejia and Daniela Vega acknowledge the support of UniAndes-GoogleDeepMind Scholarships 2022 and 2024 respectively. This work was partially supported by Azure sponsorship credits granted by Microsoft’s AI for Good Research Lab.

### v2

@@ 5.2 Gene Prediction Benchmark
Fig. 6 shows the performance of all methods for every dataset when trained under our two scenarios (raw and SpaCKLE-completed data). It is clear that the prediction performance significantly improves when applying SpaCKLE to every dataset and, in some cases, the best PCC increases to 0.36 points (AHSCC). This result pinpoints the importance of acknowledging missing data for the prediction task and proves the significance of including gene completion in ST pipelines.
@@
Comparing datasets’ difficulty, we find that the most challenging dataset to predict was MMBR (PCC=0.16), while 10XGMBSP emerged as the least difficult (PCC=0.74). When inspecting each dataset’s characteristics, we observe that the organism does not appear to have a significant impact on the difficulty of the task, as the mean prediction PCC achieved for mouse and human datasets is very similar. Furthermore, a larger number of available genes (due to better quality) facilitates prediction, which is evident in a higher average and maximum performance on the datasets with 128 genes compared to those with 32 genes. Finally, results also demonstrate that generalizing in an intra-subject manner typically makes the prediction easier than inter-subject (See Fig.7. a).
@@
We analyze the prediction performance across various tissue types. Fig. 7. b presents the performance of each dataset in SpaRED, categorized by tissue type, where the bars indicate the average PCC for each type of tissue. On average, the best prediction performance was observed for skin tissue, while the lowest was for kidney tissue. However, the distribution of tissue types in SpaRED is highly imbalanced, with some tissue types being underrepresented. Notably, both skin and kidney tissues are represented by only a single dataset, making it unreliable to draw definitive conclusions about whether certain tissues are inherently easier to predict than others. The observed differences may be influenced by dataset-specific characteristics rather than general tissue properties.
@@
[…]
@@
[…]
@@ Acknowledgments
Gabriel M. Mejia and Daniela Vega acknowledge the support of UniAndes-GoogleDeepMind Scholarships 2022 and 2024 respectively. This work was supported by Azure sponsorship credits granted by Microsoft’s AI for Good Research Lab.